OpenStudy (anonymous):

A = a+b divided by 2 when A = 93 and a= 88

OpenStudy (anonymous):

b=10

OpenStudy (lgbasallote):

arns...93-88 = 5...just saying :)

OpenStudy (anonymous):

WOOOOPS..

OpenStudy (anonymous):

o basically, if you do not understand how u got that, this is how u do it. We want to get b alone. So we have to subtract 88 by each side. That is. 5=b/2 Now multiply by 2 10=b